if the length of the hypotenuse of a right triangle is 10, find the lengths of the other sides when the area is a minimum


1
Expert's answer
2021-05-13T03:24:30-0400

A right triangle has one side equal to x, its hypotenuse equals 10. Then its third side is equal to "\sqrt{100-x^2}" .


Area of this triangle is "A(x)=\frac{x\sqrt{100-x^2}}{2}" .


Let is find minimum of the function "A(x)" on interval (0;10) (the length of the side is a positive number and to can’t be greater than length of hypotenuse)

"A^\prime (x)=\left(

\frac{x\sqrt{100-x^2}}{2}

\right)^\prime = \frac{\sqrt{100-x^2}}{2}-\frac{x^2}{2\sqrt{100-x^2}}=\frac{100-2x^2}{2\sqrt{100-x^2}}"

We put "A^\prime (x)=0"

"100-2x^2=0"

"x^2=50"

"x=5\sqrt{2}"

We compute "A(x)" at the critical point: "A(5\sqrt{2})=25"

Now we find "A(x)" at the boundaries of the interval: "A(0)=A(10)=0" .

Therefore, "x=5\sqrt{2}" is the point of maximum and "A(x)" is minimum at the boundaries, but when "x=0" or "x=10" there is no triangle.

So, there is no minimum area of this triangle.